A bank offers both adjustable and fixed rate mortgage loans on residential properties, which are classified into three categories: single-family homes, condominiums, and multi-family homes. Each loan made in 2010 was classified according to type of mortgage and type of property, as shown in the given table.

Single Family Condo Multi Family

Adjustable 1500 788 337

Fixed Rate 375 377 373

Answer the following: (Express answers as fractions or to 4 decimal places) a) What is the probability that a randomly chosen loan was for a condo? b) What is the probability that a randomly chosen loan was for a fixed rate, given that is was on a single family home? c) What is the probability that a randomly chosen loan was for a multi family home, given that it was an adjustable rate loan?
Answer #1

a)probability that a randomly chosen loan was for a condo =1165/3750 =0.3107

b) probability that a randomly chosen loan was for a fixed rate, given that is was on a single family home

=375/1875=0.2

c)

probability that a randomly chosen loan was for a multi family home, given that it was an adjustable rate loan=337/2625=0.1284